For one mole of gas the average kinetic energy is given as E. The Urms of Gas is

Grade:11

1 Answers

Arun
6 years ago
we know that
Urms = \sqrt(3RT/M)        …......(i)
 
and we also know that 
E = (3/2)RT            …............(ii)
hence from (i) & (ii) – 
Urms = \sqrt(2E/M)
 
hope it helps